Find values for m and b in the following system so that the solution to the system
is (-3, 4).
5x +7y = b
mx + y = 22

Answer by malakumar_kos@yahoo.com(315)


Find values for m and b in the following system so that the solution to the system
is (-3, 4).
5x +7y = b
mx + y = 22

solution is (-3,4) hence 5(-3)+7(4) = b
-15 +28 = b
b = 13
(-3,4) hence m(-3)+4 = 22
-3m = 22-4
-3m = 18
m = -18/3 = -6
Therefore the values of m = -6 and b = 13